OpenStudy (anonymous):

Explain, in complete sentences, how you would use the elimination method to solve the following system of equations. Part 2: Provide the solution to the system. 5x – 3y = 1 7x – 4y = 2

hero (hero):

I can help you with this

OpenStudy (anonymous):

Ok

hero (hero):

1. Solve both equations for y: \[y = \frac{1-5x}{-3}\] \[y = \frac{2-7x}{-4}\] Next, set y = y: \[\frac{1-5x}{-3} = \frac{2-7x}{-4}\]

hero (hero):

Cross Multiply, then isolate x

hero (hero):

-4(1-5x)=-3(2-7x) -4+20x = -6+21x x = 2

hero (hero):

y = 3

hero (hero):

I think that's good enough. If not, let me know

hero (hero):

I'm guessing you didn't like my steps

hero (hero):

Well at least somebody liked my steps.